# IDs for AI Systems

### v1

@@ Abstract
AI systems are increasingly pervasive, yet information needed to decide whether and how to engage with them may not exist or be accessible. A user may not be able to verify whether a system satisfies certain safety standards. An investigator may not know whom to investigate when a system causes an incident. A platform may find it difficult to penalize repeated negative interactions with the same system. Across a number of domains, IDs address analogous problems by identifying particular entities (e.g., a particular Boeing 747) and providing information about other entities of the same class (e.g., some or all Boeing 747s). We propose a framework in which IDs are ascribed to instances of AI systems (e.g., a particular chat session with Claude 3), and associated information is accessible to parties seeking to interact with that system. We characterize IDs for AI systems, argue that there could be significant demand for IDs from key actors, analyze how those actors could incentivize ID adoption, explore potential implementations of our framework, and highlight limitations and risks. IDs seem most warranted in high-stakes settings, where certain actors (e.g., those that enable AI systems to make financial transactions) could experiment with incentives for ID use. Deployers of AI systems could experiment with developing ID implementations. With further study, IDs could help to manage a world where AI systems pervade society.
@@ 1 Introduction
AI systems are becoming increasingly commonplace. While current systems can struggle to complete complex tasks (Mialon et al. 2023; Liu et al. 2023; Kinniment et al. 2023; Xie et al. 2024; Jimenez et al. 2024), capabilities seem likely to improve (Hoffmann et al. 2022; Epoch 2023; Erdil and Besiroglu 2023; Ho et al. 2024), and future AI agents could carry out a broad range of tasks with only minimal human intervention (Chan et al. 2023; Shavit et al. 2023; Chan et al. 2024). Several commercially deployed AI systems can already search the web, send emails, and write code (OpenAI 2024b; Anthropic 2024a). Even when they do not function reliably, AI systems might still be widely used, whether because of cost advantages, hype, or the externalization of their harms (De La Garza 2020; BBC 2021; Raji et al. 2022).
@@
At the same time, information to make decisions about engaging with AI systems may not exist or be accessible. Although there may be obligations to inform parties that they are interacting with an AI system (e.g., EU AI Act Article 50.1 (Parliament 2024)), those parties may not know with which AI system they are interacting. For example, a party that knowingly interacts with an AI system may not be aware that their system is relatively more vulnerable to adversarial attacks (Zhan et al. 2024). Awareness of such elevated risk could justify additional precautions, such as reviewing the system’s actions or abstaining from interaction altogether. Furthermore, the same party—or an investigator—may lack the information to pursue recourse if the system causes harm. Information about the system or the interaction (e.g., whether a system behaved according to safety standards, or the identity of the deployer) could aid incident investigation, allocation of liability, or other legal action (Buiten, de Streel, and Peitz 2023; Buiten 2024; Kolt 2024; Wills 2024).
@@
Across numerous domains, IDs held by software, assets, individuals, and organizations address analogous problems. IDs can help to ascertain compliance with standards or regulation. For instance, an individual aircraft’s tail number is associated to the aircraft’s incident and maintenance history, which could inform safety assessments from regulators or aircraft operators. Furthermore, IDs can help to establish whether trust is warranted. A website’s valid HTTPS certificate assures users of the website domain’s authenticity, and provides a way for users to establish a secure communication channel. [^1] Despite a lack of centralized enforcement, HTTPS gradually became the norm due to widespread awareness of its security benefits and collective advocacy from web browsers, search engines, and other organizations (Encrypt 2024; Hancock 2021). Finally, IDs can facilitate redress. Serial numbers on consumer products enable customer support, product recalls, and attribution of liability. A key feature that enables all of the above functions is that IDs are specific to particular entities (e.g., a particular Boeing 747), although they may also contain information about other entities of the same class (e.g., some or all Boeing 747s).
@@
Identifying particular AI systems, which we term instances, could be similarly useful. An instance corresponds to a context window and a user. [^2] This definition of instance does not take memory into account and is anchored upon current language models. While it will suffice for this work, see Appendix A for a more general definition. For example, a particular user’s chat session with ChatGPT (with e.g. a GPT-4 backend) is an instance, and that instance could interact with other parties by sending an email through the user’s account (OpenAI 2024b). Separate instances can behave differently, whether because of user instructions (Shanahan, McDonell, and Reynolds 2023; Wei et al. 2023; Bai et al. 2022; Zou et al. 2023; Agarwal et al. 2024), hijacking by an attacker (Greshake et al. 2023; Zhan et al. 2024), or malfunction. As such, instance-specific information could aid decisions about interactions. For example, repeated negative interactions with the same instance may cause a party to reject further interaction from the instance. The ability to distinguish instances from one another could aid incident investigation (e.g., a client files a complaint about an instance) and allocation of liability (e.g., if a third party hijacked the instance). Moreover, instance-level identification would allow flexible grouping of instances. For example, an incident investigator could be interested in all the instances that interacted with a particular party.
@@
Yet, current mechanisms do not identify instances. System documentation (Gebru et al. 2021; Mitchell et al. 2019; Gilbert et al. 2023; Bommasani et al. 2023) provides information about systems, but such documentation is not an ID for instances or systems (however, an ID could include such documentation). API tokens for services, such as for hotel booking (OpenAI 2023a), do identify entities that use the tokens, but are usually only user- or device-specific. Even if API requests do include a string with the name of the system, any actor could mimic the string in response to the same service. Finally, user accounts only separate the activities of AI systems from different users.
@@
To prepare for a world with ubiquitous AI interactions, we propose (instance-level) IDs for AI systems. An AI ID is a container for 1) an identifier and 2) attributes. An identifier is a unique string that refers to an instance. [^3] It may be desirable in the future to have identifiers that refer to even more granular parts of an AI system. For example, one could have an identifier refer to an instance as it operated between times $t$ and $t+10$. An identifier could be randomly generated, or could itself encode some information about the instance (e.g., the identifier could include the exact time at which the instance started operating). An attribute is any information that could pertain to the instance, and could include behaviour, properties, context, or relationships to other instances or systems. An identifier enables the association of attributes to an instance, similar to how serial numbers associate information to a particular product. Attributes could be specific to an instance (e.g., prior incidents associated with the instance), or could apply more broadly to other instances (e.g., a system [^4] A system (e.g., ChatGPT) could use different models (e.g., GPT-4, GPT-3.5) as a backend. We consider system cards to include model cards (Mitchell et al. 2019). card). An ID could directly include attributes (e.g., the name of the deployer) or could link to them (e.g., a link to a database of prior incidents).
@@
Different instances or systems could warrant attributes of differing levels of granularity or detail. For example, attaching a user identifier to an ID may only be appropriate in high-stakes settings, such as when a company uses an AI system to interface with critical infrastructure. Prior incidents associated with an instance would likely be more useful for future AI agents (Chan et al. 2023; Shavit et al. 2023; Chan et al. 2024; Gabriel et al. 2024), which could act autonomously and persist over long durations.
@@
Since IDs would contain information about AI systems, they could complement measures that attempt to verify whether AI systems—or their artifacts—are present. Watermarks (Liu et al. 2024; Wang et al. 2021) embed origin information in AI outputs, while content provenance measures (C2PA 2023) embed such information in metadata. Both types of techniques could embed IDs. Other measures, like CAPTCHAs (Shet 2014), verify that a human is performing an action, so as to reduce service abuse. If AI systems become essential for many tasks, such as web search and account registration, it may be useful to allow them to bypass CAPTCHAs in exchange for presenting an ID. [^5] Advances in AI capabilities may render CAPTCHAs ineffective. We discuss this possibility further in Section 4. Parties could use this ID to track and disincentivize abuse.
@@ 1.1 Contributions
We propose IDs for AI systems. First, we characterize the central properties of IDs. Second, we argue that there will likely be demand for IDs from several key actors, especially in high-stakes settings. We also explore potential ways for these actors to incentivize ID use. Third, we analyze potential technical implementations of IDs. Finally, we investigate some of the limitations of IDs and of our analysis, including privacy and security risks of IDs, as well as uncertainty about the broader societal consequences of IDs for AI systems.
We recommend limited experimentation with IDs. IDs seem most warranted in high-stakes settings, where certain actors could experiment with incentives for ID use. For example, actors that provide interfaces for AI systems to carry out financial transactions could impose rate limiting whenever IDs are not present. However, instances without IDs should still be allowed to access services. Deployers of AI systems could experiment with developing the technical infrastructure to enable IDs. ID implementation in decentralized deployments will require more study.
Figure 1: IDs contain a unique identifier along with attributes (e.g., a system card or [a link to] previous incidents). We also display some potential actions that parties might take based on information in an ID.
@@ 2 Definitions
A service is software that allows an entity to perform tasks. For example, software which an individual uses to perform online banking is a service. An (AI) plugin (OpenAI 2023a; Richards 2023) is software that allows AI systems to interact with services. For instance, plug-ins allow GPT-4 to interact with web search, Wikipedia, and Twitter (Significant-Gravitas 2024). A plugin developer is not necessarily the same as a service provider—the actor that develops and maintains the service—since the former can take advantage of existing interfaces for services. Third parties (plugin developer) use Microsoft’s (service provider) existing software interface for Bing to allow GPT-4 to perform web searches (Richards 2023).
@@ 3 Characterizing IDs
An ID is a container (Korenhof et al. 2014) for 1) an identifier (that corresponds to an instance) and 2) attributes. We characterize a design space for IDs: the attributes it includes, to whom it is accessible, and to what extent it is verifiable. Our goal in this section is to investigate possible designs and why they may be desirable, rather than to prescribe specific choices.
@@ 3.1 Attributes
  • Behaviour: prior incidents (Wei and Heim 2024)
  • Properties: information in a system card (Mitchell et al. 2019; Gebru et al. 2021); the results of evaluations (Shevlane et al. 2023; Weidinger et al. 2023)
  • Context: the system prompt; external memory (Wang et al. 2023)
@@
Behaviour and properties could straightforwardly inform AI interaction decisions. Context could inform parties about potential behaviour (e.g., there is a jailbreak in the system prompt or external memory (Cohen, Bitton, and Nassi 2024)). Lastly, relationship information could aid incident investigation and response. Suppose a user instructs an instance to carry out a personalized influence campaign. The instance could create many descendant instances, each of which could target an individual person. If the ID of a descendant instance is linked to the ID of the original, ancestor instance, it could be easier to investigate and resolve such misuse. In Sections 5 and 6, we discuss when this ID linking is possible.
@@
3.2 Accessibility
The accessibility of an ID refers to who can view it and the conditions under which they can do so.
Parties that can view the ID: We distinguish between primary parties and secondary parties. Primary parties receive an ID when interacting with the corresponding instance. Secondary parties receive the ID through other means, such as directly from a primary party, or from records. Some examples include:
@@ 3.2 Accessibility
Making ID accessible to primary parties corresponds to attaching IDs to certain outputs from the instance. Some primary parties and how IDs may be attached include:
@@
  • If user identification (e.g., from a know-your-customer process (Egan and Heim 2023)) is included in an ID, it may only be appropriate for government authorities to access this identification for e.g. regulatory purposes
@@
Actors implementing IDs could selectively hide information from different primary parties. Those same actors could also process attributes in a privacy-preserving way (Trask et al. 2023; Sporny et al. 2024) so as to reveal only information that is relevant to the party interacting with an instance. Even so, preventing transmission of information to other parties may be difficult. For example, once an actor obtains user identification, it could—intentionally or not—transfer that information to other actors.
Persistence: IDs could be accessible for varying amounts of time after an instance has ceased to operate. Practical considerations, such as storage capacity, could impose limitations. The appropriate duration could also depend upon the application domain. Analogously, financial institutions have obligations to maintain records for set periods of time (noa 2022).
@@ 3.3 Verifiability
Suppose an author creates and sends an ID to a party that is interacting, or seeking to interact, with an instance. For the party to trust that the ID from the author provides accurate information about the instance, the following are necessary (but not sufficient; see further discussion below).
@@
    The party’s received ID is the same as the ID that the author sent
@@
We consider IDs (i.e., the ID system) to be verifiable if they defend against these threats. To defend against tampering, it should be easy to check if an attacker has modified the ID. To combat ID spoofing, it should be computationally difficult for the attacker to spoof authorship. For instance spoofing, it should be easy to check whether an ID indeed corresponds to a particular instance. In Sections 5 and 6, we analyze potential implementations that attempt to counteract these threats. See also (Sporny et al. 2022, 2024; Microsoft 2023), which are increasingly popular standards to implement such ID requirements in a domain-agnostic way.

### v3

@@ Abstract
AI systems are increasingly pervasive, yet information needed to decide whether and how to engage with them may not exist or be accessible. A user may not be able to verify whether a system has certain safety certifications. An investigator may not know whom to investigate when a system causes an incident. It may not be clear whom to contact to shut down a malfunctioning system. Across a number of domains, IDs address analogous problems by identifying particular entities (e.g., a particular Boeing 747) and providing information about other entities of the same class (e.g., some or all Boeing 747s). We propose a framework in which IDs are ascribed to instances of AI systems (e.g., a particular chat session with Claude 3), and associated information is accessible to parties seeking to interact with that system. We characterize IDs for AI systems, provide concrete examples where IDs could be useful, argue that there could be significant demand for IDs from key actors, analyze how those actors could incentivize ID adoption, explore a potential implementation of our framework for deployers of AI systems, and highlight limitations and risks. IDs seem most warranted in settings where AI systems could have a large impact upon the world, such as in making financial transactions or contacting real humans. With further study, IDs could help to manage a world where AI systems pervade society.
@@ 1 Introduction
AI systems are becoming increasingly commonplace. While current systems can struggle to complete complex tasks (Mialon et al., 2023; Liu et al., 2023; Kinniment et al., 2023; Xie et al., 2024; Jimenez et al., 2024), capabilities seem likely to improve (Hoffmann et al., 2022; Epoch, 2023; Erdil & Besiroglu, 2023; Ho et al., 2024), and future AI agents could carry out a broad range of tasks with only minimal human intervention (Chan et al., 2023; Shavit et al., 2023; Chan et al., 2024). Several commercially deployed AI systems can already search the web, send emails, and write code (OpenAI, 2024b; Anthropic, 2024a). Even when they do not function reliably, AI systems might still be widely used, whether because of cost advantages, hype, or the externalization of their harms (De La Garza, 2020; BBC, 2021; Raji et al., 2022).
@@
At the same time, information to make decisions about engaging with AI systems may not exist or be accessible. Although there may be obligations to inform parties that they are interacting with an AI system (e.g., EU AI Act Article 50.1 (Parliament, 2024)), those parties may not know with which AI system they are interacting. For example, a party that knowingly interacts with an AI system may not be aware that their system is relatively more vulnerable to adversarial attacks (Zhan et al., 2024). Awareness of such elevated risk could justify additional precautions, such as reviewing the system’s actions or abstaining from interaction altogether. Furthermore, the same party—or an investigator—may lack the information to pursue recourse if the system causes harm. Information about the system or the interaction (e.g., whether a system behaved according to safety standards, or the identity of the deployer) could aid incident investigation, allocation of liability, or other legal action (Buiten et al., 2023; Buiten, 2024; Kolt, 2024; Wills, 2024).
@@
Across numerous domains, IDs held by software, assets, individuals, and organizations address analogous problems. IDs can help to ascertain compliance with standards or regulation. For instance, an individual aircraft’s tail number is associated to the aircraft’s incident and maintenance history, which could inform safety assessments from regulators or aircraft operators. Furthermore, IDs can help to establish whether trust is warranted. A website’s valid HTTPS certificate assures users of the website domain’s authenticity, and provides a way for users to establish a secure communication channel. [^1] Despite a lack of centralized enforcement, HTTPS gradually became the norm due to widespread awareness of its security benefits and collective advocacy from web browsers, search engines, and other organizations (Encrypt, 2024; Hancock, 2021). Finally, IDs can facilitate redress. Serial numbers on consumer products enable customer support, product recalls, and attribution of liability. A key feature that enables all of the above functions is that IDs are specific to particular entities (e.g., a particular Boeing 747), although they may also contain information about other entities of the same class (e.g., some or all Boeing 747s).
@@
Identifying particular AI systems, which we term instances, could be similarly useful. Given their increasing prevalence, we focus here on digital-assistant-like systems that that respond to user direction, usually (but not necessarily) through natural language (Gabriel et al., 2024). In this context, an instance corresponds to a user and a history of interaction. [^2] This definition of instance does not take memory into account. While it will suffice for this work, see Appendix A for a more general definition. For example, a particular user’s chat session with ChatGPT (with e.g. a GPT-4 backend) is an instance. Separate instances can behave differently, whether because of user instructions (Shanahan et al., 2023; Wei et al., 2023; Bai et al., 2022; Zou et al., 2023; Agarwal et al., 2024), hijacking by an attacker (Greshake et al., 2023; Zhan et al., 2024), or malfunction. As such, instance-specific information could aid decisions about interactions. For example, the ability to distinguish instances from one another could aid incident investigation (e.g., a client files a complaint about an instance) and allocation of liability (e.g., if a third party hijacked the instance).
@@
Yet, current mechanisms are not granular enough to identify instances. System documentation (Gebru et al., 2021; Mitchell et al., 2019; Gilbert et al., 2023; Bommasani et al., 2023) provides information about systems, but such documentation is not an ID for instances or systems (however, an ID could include such documentation). API tokens for services, such as for hotel booking (OpenAI, 2023a), do identify entities that use the tokens, but are usually only user- or device-specific, and do not allow identification of an entity across different services. Even if API requests do include a string with the AI system’s name, any attacker could mimic it. Finally, user accounts only separate the activities of AI systems from different users. We summarize these differences and discuss further identity systems in table 1.
Existing Systems Differences with AI IDs
Proof of personhood (Borge et al., 2017) A proof of personhood identifies a unique human in an anonymous way, but does not identify AI systems.
ORCID An ORCID is meant to point to a particular researcher, but there is no reliable way to verify an entity’s claim that a particular ORCID corresponds to it.
DOI A DOI is meant to provide a unique, persistent identifier of a digital object, but there is no reliable way to verify an entity’s claim that a particular DOI corresponds to it.
Watermarks Watermarks are meant to indicate AI provenance in content and do not contain information about particular AI systems.
C2PA (C2PA, 2023) C2PA is a standard for making verifiable claims about the provenance of content. Such claims can point to an ID, but are not IDs in themselves.
CAPTCHAs CAPTCHAs are meant to detect presence of non-humans, rather than identify them.
API tokens API tokens are user- or device-specific and do not identify entities across different APIs.
User accounts User accounts only separate the activities of AI systems from different users.
System documentation, such as a system card System documentation provides information about systems, but is not an ID for instances or systems.
Table 1: We collate a number of existing identity-like systems, some of which do not apply to AI systems, and describe their differences with AI IDs.
@@
To prepare for a world with ubiquitous AI interactions, we propose (instance-level) IDs for AI systems. An AI ID is a container for 1) an identifier and 2) attributes. An identifier is a unique string that refers to an instance. [^3] It may be desirable in the future to have identifiers that refer to even more granular parts of an AI system. For example, one could have an identifier refer to an instance as it operated between times $t$ and $t+10$. An identifier could be randomly generated, or could itself encode some information about the instance (e.g., the identifier could include the exact time at which the instance started operating). An attribute is any information that could pertain to the instance, and could include behaviour, properties, context, or relationships to other instances or systems. An identifier enables the association of attributes to an instance, similar to how serial numbers associate information to a particular product. Attributes could be specific to an instance (e.g., prior incidents associated with the instance), or could apply more broadly to other instances (e.g., a system [^4] A system (e.g., ChatGPT) could use different models (e.g., GPT-4, GPT-3.5) as a backend. We consider system cards to include model cards (Mitchell et al., 2019). card). An ID could directly include attributes (e.g., the name of the deployer) or could link to them (e.g., a link to a database of prior incidents).
@@
Different instances or systems could warrant attributes of differing levels of granularity or detail. For example, attaching a user identifier to an ID may only be appropriate in high-stakes settings, such as when a company uses an AI system to interface with critical infrastructure. Prior incidents associated with an instance would likely be more useful for future AI agents (Chan et al., 2023; Shavit et al., 2023; Chan et al., 2024; Gabriel et al., 2024), which could act autonomously and persist over long durations.
@@
IDs would contain information about AI systems and are thus are different from measures that attempt to verify whether AI systems—or their artifacts—are present. Watermarks (Liu et al., 2024; Wang et al., 2021) embed origin information in AI outputs, while content provenance measures (C2PA, 2023) embed such information in metadata. Both types of techniques could embed IDs. Other measures, like CAPTCHAs (Shet, 2014), verify that a human is performing an action, so as to reduce service abuse. If AI systems become essential for many tasks, such as web search and account registration, it may be useful to allow them to bypass CAPTCHAs in exchange for presenting an ID. [^5] Advances in AI capabilities may render CAPTCHAs ineffective. We discuss this possibility further in Section 5. Parties could use this ID to track and disincentivize abuse.
@@ 1.1 Contributions
We propose IDs for AI systems. First, we characterize the central properties of IDs. Second, we provide concrete examples where IDs could be useful. Third, we argue that there will likely be demand for IDs from several key actors, especially in high-stakes settings. We also explore potential ways for these actors to incentivize ID use. Fourth, we analyze a potential implementation of IDs. Finally, we investigate some of the limitations of IDs and of our analysis, including privacy and security risks of IDs, as well as uncertainty about the broader societal consequences of IDs for AI systems.
We recommend limited experimentation with IDs. IDs seem most warranted in settings where AI systems could have a large impact upon the world, such as in making financial transactions or contacting real humans. For example, actors that provide interfaces for AI systems to carry out financial transactions could impose rate limiting whenever IDs are not present. However, instances without IDs should still be allowed to access services. Deployers of AI systems could experiment ID implementation.
Figure 1: IDs contain a unique identifier along with attributes (e.g., a system card, certifications, or a link to previous incidents). We also display some potential actions that parties might take based on information in an ID.
@@ 2 Definitions
A service is software that allows an entity to perform tasks. For example, software which an individual uses to perform online banking is a service. An (AI) plugin (OpenAI, 2023a; Richards, 2023) is software that allows AI systems to interact with services. For instance, plugins allow GPT-4 to interact with web search, Wikipedia, and Twitter (Significant-Gravitas, 2024). A plugin developer is not necessarily the same as a service provider—the actor that develops and maintains the service—since the former can take advantage of existing interfaces for services. Third parties (plugin developer) use Microsoft’s (service provider) existing software interface for Bing to allow GPT-4 to perform web searches (Richards, 2023).
@@ 3 Characterizing IDs
An ID is a container (Korenhof et al., 2014) for 1) an identifier (that corresponds to an instance) and 2) attributes. We characterize a design space for IDs: the attributes it includes, to whom it is accessible, and to what extent it is verifiable. Our goal in this section is to investigate possible designs and why they may be desirable, rather than to prescribe specific choices.
@@ 3.1 Attributes
  • Behaviour: prior incidents (Wei & Heim, 2024)
  • Properties: information in a system card (Mitchell et al., 2019; Gebru et al., 2021); the results of evaluations (Shevlane et al., 2023; Weidinger et al., 2023)
  • Context: the system prompt; external memory (Wang et al., 2023)
@@
Behaviour and properties could straightforwardly inform AI interaction decisions. Context could inform parties about potential behaviour (e.g., there is a jailbreak in the system prompt or external memory (Cohen et al., 2024)). Lastly, relationship information could aid incident investigation and response. Suppose a user instructs an instance to carry out a personalized influence campaign. The instance could create many descendant instances, each of which could target an individual person. If the ID of a descendant instance is linked to the ID of the original, ancestor instance, it could be easier to investigate and resolve such misuse. In Section 6, we discuss when this ID linking is possible.
@@
3.2 Access
An ID is simply an abstraction that contains or links to information. That abstraction can be presented in several potential ways, such as on a separate web page or in a pop-up window. We explore design decisions around who can access an ID and how it is presented to them.
Parties that can access the ID: We distinguish between primary parties and secondary parties. Primary parties receive an ID when interacting with the corresponding instance. Secondary parties receive the ID through other means, such as directly from a primary party, or from records. Some examples include:
@@ 3.2 Access
Presenting an ID to primary parties corresponds to linking IDs to the corresponding instance’s outputs. Content provenance standards (C2PA, 2023) can help to ensure that such linking is verifiable (see Section 3.3). Some primary parties and how IDs may be attached include:
@@
  • If user identification (e.g., from a know-your-customer process (Egan & Heim, 2023)) is included in an ID, it may only be appropriate for government authorities to access this identification for e.g. regulatory purposes
@@
Actors implementing IDs could selectively hide information from different primary parties. Those same actors could also process attributes in a privacy-preserving way (Trask et al., 2023; Sporny et al., 2024) so as to reveal only information that is relevant to the party interacting with an instance. Even so, preventing transmission of information to other parties may be difficult. For example, once an actor obtains user identification, it could—intentionally or not—transfer that information to other actors.
Persistence: IDs could be accessible for varying amounts of time after an instance has ceased to operate. Practical considerations, such as storage capacity, could impose limitations. The appropriate duration could also depend upon the application domain. Analogously, financial institutions have obligations to maintain records for set periods of time (noa, 2022).
@@ 3.3 Verifiability
Suppose an author creates an ID (or a link to an ID) for an instance. A party interacts, or seeks to interact, with the instance. The party receives an ID. For the party to trust that the ID provides accurate information about the instance, the following are necessary (but not sufficient; see further discussion below).
@@
    The party’s received ID is the same as the author’s created ID
@@
We consider IDs (i.e., an ID system) to be verifiable if they defend against these threats. We provide an illustration of these threat models in Section 3.3.
Existing technology could be drawn upon to provide verifiability. For digital-assistant-like AI systems (e.g., GPT-4), any digital communications usually would take place over TLS/SSL (Rescorla, 2018), which would address tampering. We leave tampering of other AI systems’ IDs (e.g., IDs for self-driving cars) to future work. To combat ID spoofing, authors could digitally sign the ID (potentially with public key ownership verified by a trusted third party). For instance spoofing, there are existing standards for binding a digital object, such as an ID, to a piece of content, such as an AI system’s output. For example, C2PA involves binding provenance claims to content (C2PA, 2023), and calls the binded object a manifest. Such binding could ensure that IDs are associated to particular outputs from an instance in a verifiable way. [^7] If desired, the verifiability of IDs over extended interactions could be ensured by linking manifests to each other cryptographically, such as in a blockchain. We emphasize that our examples here are meant to be illustrative, not prescriptive. Security vulnerabilities in existing technologies may weaken verifiability and require new solutions.
Figure 2: We illustrate how existing technologies can help ensure the verifiability of IDs against the threat models in Section 3.3. We emphasize that our examples of existing technologies are meant to be illustrative, not prescriptive.
@@
4 How Could IDs Help?
We provide some hypothetical case studies where IDs could be useful and compare IDs to alternatives. AI agents—systems which act autonomously to achieve goals with little to no user direction, instruction, or supervision (Chan et al., 2023; Shavit et al., 2023; Gabriel et al., 2024)—feature prominently in our scenarios.
4.1 Shutting Down Malfunctioning Agents
Scenario: Suppose a user runs an AI agent to provide continual recommendations across many platforms and services (e.g., for media, shopping etc). [^8] The idea for this scenario is from Zittrain (2024). To generate such recommendations, the agent interacts with plugins for various web services, such as search engines, online discussion fora, etc. The agent provides its ID to each service. Unbeknownst to the user, the agent begins to malfunction (e.g., by chance, prompt injection) and disrupt services (e.g., spamming discussion fora, sending malformed requests to services or overloading them, conducting prompt injection attacks on other AI agents). Services interacting with the agent flag its malfunction. Service providers use information in the ID to notify the deployer, and block further interaction with the agent. The deployer receives multiple reports of malfunction for the same ID and investigates the agent, potentially shutting it down and notifying the user.
Alternatives to IDs: To block further interaction with the agent, a service provider can simply revoke its API token and potentially block the requesting IP address. These actions would not prevent the agent from accessing new services, however. As well, API tokens would not identify an agent across different services to enable shutdown if necessary. Alternatively, services could notify the user of agent malfunction. With a good user interface, users could notice such malfunction and shut down their agent. If users are careless or rarely supervise their agent (perhaps they have many such agents running in the background), it seems useful to have IDs as a backup option.
Limitations of the scenario: It is unclear how impactful the disruption of services would be. Deployers could also be unresponsive to notifications from service providers.
4.2 Verifying Certification
Scenario: A user wishes to interact with an agent from another person. The user wants to make sure that the agent does not act in an undesirable way. For example, the user would want to make sure that a malicious third party did not prompt inject (Greshake et al., 2023) the agent (perhaps because the user is passing sensitive information or is accessing resources the agent will send over). A trusted third party, such as a deployer or an auditor, checks the interaction history of the agent at runtime for prompt injections (e.g., with a tool like Beurer-Kellner et al. (2024)). After verifying the absence of prompt injections, the trusted third party adds a certification onto the agent’s ID. The user sees the certification on the ID and continues interaction.
Alternatives to IDs: The trusted third party could send the certification directly to a user requesting it, but some sort of identifier would still be needed to ensure that the certification corresponds to the agent with which the user is interacting.
Limitations of the scenario: Certification would raise privacy concerns because the trusted third party would be able to see the whole interaction history. Privacy-preserving ways to perform such audits would be valuable to explore (Trask et al., 2023; 2024). A potential alternative would be for the auditor to run behavioural tests on the agent at runtime, and link the results of the tests to the ID.
4.3 Investigating Scam Calls
Scenario: Suppose a user spins up an AI agent to carry out scam phone calls. The agent creates multiple sub-agents—potentially using different AI deployers—and directs each one to carry out scam calls through a telephone service. The service receives each sub-agent’s ID, each of which contains information about the deployer and a connection to the original agent’s ancestor ID. The service provider connects the sub-agent to a telephone network and creates a call detail record (CDR), linking the call with the sub-agent’s ID. Some scam victims notify the operator of the telephone service. The telephone provider notifies law enforcement, who investigates the relevant CDRs and IDs. Law enforcement notifies the deployers of the sub-agents, potentially after obtaining a warrant so as to obtain access to the deployers’ logs. Through ID linkages between parent and child agents, law enforcement and deployers discover the agent and user behind the spam calls.
Alternatives to IDs: The main alternative would be for law enforcement to ask ISPs to perform IP address tracking, if the telephone provider records the IP addresses of the sub-agents. Interacting with ISPs (and potentially obtaining a warrant) could take additional time and effort. Conservatively, in this case IDs could help to save time during incident investigation, but not necessarily enable something that was otherwise impossible.
Limitations of the scenario: The main limitation of this scenario is the provision of IDs from the sub-agents to the service provider. Users could use deployers that do not implement IDs, or could run their own agents to avoid passing IDs to APIs. A potential response is for the telephone service provider to require IDs before provisioning telephone access.